# 111preprint

### v1

@@ Abstract.
Our goal is to model and experimentally assess trust evolution to predict future beliefs and behaviors of human-robot teams in dynamic environments. Research suggests that maintaining an appropriate level of trust among team members in a human-robot team is vital for successful team performance. Indeed, research suggests that trust is a multi-dimensional and a latent entity that relates to past experiences and future actions in a complex manner. Employing a human-robot collaborative task, we design an optimal assistance-seeking strategy for the robot using a Partially Observable Markov Decision Process (POMDP) framework. In the task, the human partner supervises an autonomous mobile manipulator collecting objects in an environment. The supervisor’s responsibility is to ensure that the robot safely executes its task. The robot can either choose to attempt to autonomously collect the object or seek human assistance. The human supervisor actively monitors the robot’s activities, offering assistance upon request, and intervening if they perceive that the robot may fail. In this setting, human trust is the hidden state, and the primary objective is to optimize team performance. We execute two sets of human-robot interaction experiments. The data from the first experiment are used to estimate POMDP parameters, which are used to compute an optimal assistance-seeking policy evaluated in the second experiment. The estimated POMDP parameters reveal that, for most participants, human intervention is more probable when trust is low, particularly in high-complexity tasks. Notably, our estimates suggest that the robot’s action of asking for human assistance in high-complexity tasks can positively impact human trust. Our experimental results show that the proposed trust-aware policy results in a better performance as compared to an optimal trust-agnostic policy. By comparing model estimates of human trust, obtained using only behavioral data, with the collected self-reported trust values, we show that model estimates are isomorphic to self-reported responses.
isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Autonomous systems are becoming increasingly prevalent and are being deployed and used in various sectors such as healthcare, agriculture, military, and transportation. Despite the increasing capabilities of autonomy, these systems still require human supervision and assistance when working in complex and uncertain environments(Selma2017, ; akash2020human, ; peterssupervisor2015, ; gupta2024structural, ). However, the increased capabilities of the robots are leading to novel landscapes of human-robot interaction (HRI) wherein the robots are deployed in roles requiring complex social behavior with human partners.
@@
The classical approach to studying the impact of robot reliability, transparency, and human workload on human trust treats trust as a static parameter (hoff2015trust, ). However, recent studies have focused on understanding the dynamic evolution of human-automation trust, focusing on factors such as the quality of robot performance, transparency of the robot’s operation, and the human agent’s attitude and characteristics, such as forgetfulness and self-confidence, on the trust dynamics (lee1992trust, ; lee1994trust, ; manzey2012human, ; desai2013impact, ; akash2017dynamic, ; yang2017evaluating, ; Wang2015DynamicRS, ; chen2020trust, ; akash2020human, ; williams2023computational, ). Hancock et. al. (hancock2011meta, ) present a comprehensive review of various factors affecting human trust, and categorize them into three major groups: robot-related, human-related, and environment-related factors. They discuss that human trust is most influenced by robot-related factors, which include robot reliability, performance, design, and abilities. This is followed by the environment-related factors which include task complexity, type, and physical environment. Among the three main categories, human-related factors such as demographics, personality, expertise, and personality have the least influence on trust in human-robot interaction.
@@
Deterministic linear models use a linear time-invariant dynamical system, where the states are the current trust level, cumulative trust level, and expectation bias (akash2017dynamic, ; yang2017evaluating, ; Wang2015DynamicRS, ). The input to these linear models includes human experience, which is determined by the performance and reliability of the robot. Variants of these models have been introduced to capture mutual trust between humans and robots (rahman2018mutual, ; Wang2014, ). These models have been used in several contexts, including, exploring the impact of information transparency in reconnaissance missions (yang2017evaluating, ), investigating human reliance on driving assistance systems (akash2017dynamic, ), examining subtask allocation in collaborative assembly tasks (rahman2018mutual, ), and addressing scheduling challenges in teleoperation for underwater robotic navigation (Wang2014, ).
Stochastic linear models for the evolution of trust (azevedo2021real, ) use linear time-invariant dynamical systems and a linear measurement model with Gaussian process and measurement noises. Besides stochastic linear models, many of the stochastic trust models treat the trust as a discrete (usually binary) hidden state and estimate its distribution conditioned on variables such as robot performance and human actions within a Bayesian framework(OPTIMO, ; liao2024trust, ). A subclass of these models are POMDP-based models (chen2020trust, ; zahedi2023trust, ; akash2020human, ; 10.5555/2906831.2906852, ; DBLP:journals/corr/abs-2009-11890, ). These models treat human mental states, such as trust and workload, as discrete hidden variables. The state transition can depend on robot actions, human actions, and contextual information such as environmental state and task outcome. These models have been applied to design task schedules in human-supervised robotic operations, minimizing human interventions (chen2020trust, ) and monitoring rates (zahedi2023trust, ). Additionally, they have been employed to develop optimal recommendations and associated levels of explanation based on human trust and workload (akash2020human, ; 10.5555/2906831.2906852, ). In this work, we use a POMDP-based stochastic trust model. In contrast to existing works, our focus is on a scenario where the robot can seek assistance from a human operator, preventing a loss of trust and possibly repairing it, and design an optimal assistance-seeking policy that can lead to better performance. In addition to the models referenced above, refer to (wang2023human, ) for a more detailed description of other developed trust models.
@@
In addition to a dynamic trust model, real-time estimation of trust is vital for the design of trust-based policies that are robust to modeling uncertainties. Traditionally, human trust is assessed through self-reporting, such as surveys (schaefer2016measuring, ; lee1992trust, ; muir1989operators, ). However, a limitation of this approach is its impracticality for frequent trust survey completion during the experiment. Psycho-physiological measurements(gebru2022review, ) have also been used to infer the trust state; however, the sensing devices required to collect these measurements, e.g., an EEG head cap, limit their applicability. For a more detailed review of trust measurements, refer to (gebru2022review, ). These challenges have prompted the researchers to rely purely on context and behavioral data, such as human actions, to estimate trust. This approach has been implemented and tested in several contexts such as reconnaissance mission (akash2020human, ), autonomous driving (azevedo2021real, ; DBLP:journals/corr/abs-2009-11890, ), and supervision(OPTIMO, ). We focus on a similar methodology for estimating human trust: subjective measures and behavioral data.
@@
In this paper, in the context of human-supervised object collection tasks, we design and validate trust-aware policies for an autonomous system seeking assistance from a human supervisor. The major contributions of this work are the following. First, we develop a novel experiment to explore how a robot’s requests for assistance influence human trust. Second, we use an Input-Output Hidden Markov Model (IOHMM) (536317, ) to capture the evolution of the hidden human trust state. In our formulation, the inputs are the robot’s action, environmental complexity, and the outcome or experience in the previous trial, characterized by the success or failure of the robot in accomplishing its goals, and the output is the human action. The present research sought to develop a model of trust in human-robot teams. It differed from previous modeling efforts in 3 different ways. First we investigated trust evolution in a human autonomous collaborative task where we evaluated the human’s intervention in the robot’s actions and we sought to understand how human trust is affected if the robot asked for assistance. Using data from multiple participants across several trials, we estimate the parameters of an IOHMM model to explore how assistance-seeking, environmental complexity and prior experience influence human trust and action. The trust estimates with the estimated IOHMM parameters are compared with self-reported survey data We then employ a POMDP framework to design a trust-aware assistance strategy, comparing its effectiveness with a baseline trust-agnostic policy in further experiments.
@@
The paper is structured as follows. Section 2 defines the collaborative object collection task. Section 3 reviews key ideas of IOHMM and POMDP and discusses the design of the optimal trust-aware assistance-seeking policy. The estimated human behavioral model and its implications are described in Section 4. Using the human-in-the-loop experiment, the trust-aware and trust-agnostic policies are compared in Section 5, Section 6 provides an elaborated discussion related to our results and we conclude in Section 7.
@@ 2. Human-Supervised Robotic Object Collection Experiment
Figure 1. Experimental Setup. The top figure shows the world view of the environment available to the human supervisor. The bottom left and middle figures show the local view and end-effector view used by humans during teleoperation. The bottom right figure shows the operation status. The setup uses ROS-Gazebo (Gazebo, ) and resources available in (downs2022google, ; ROSMoMa, ).
@@
The objects to be collected are placed in two different types of environments: low complexity $C^{L}$ or high complexity $C^{H}$. The complexity of the environment is determined based on the presence of obstacles, obstructing the direct path between the manipulator and the object to be collected. In our experiment, the complexity of trials, is an independent and identically distributed (i.i.d) Bernoulli random variable, with $p_{C^{H}}$ representing the probability of the complexity being high.
The outcome of each trial can either be a success or failure. Specifically, autonomous object collection is classified as a failure if the robot collides with its surroundings and/or fails to deposit the object in the bin safely. On the other hand, an autonomous collection is deemed a success if the robot safely collects and deposits the object. In our experiment, the success of autonomous collection in low complexity and high complexity, are mutually independent i.i.d Bernoulli variables with success probabilities $p_{L}^{\textup{suc}}$ and $p_{H}^{\textup{suc}}$ respectively.
@@ 3. Trust-aware Optimal Assistance-seeking
Figure 2. Collaborative object collection setup. The human behavioral model with trust as a hidden state yields the likelihood of human intervention and is used by the robot to design an optimal policy in a POMDP setting.
@@ 3.1. Background: IOHMM and POMDP
Consider a Markov chain with state $\{\bar{S}_{t}\in\bar{\mathcal{S}}\}_{t\in\mathbb{N}}$ with an input $\{u_{t}\in\mathcal{U}\}_{t\in\mathbb{N}}$, i.e., $\bar{S}_{t}$ is a random process such that the probability distribution of $\bar{S}_{t+1}$ is completely determined by $\bar{S}_{t}$ and $u_{t+1}$. Here, $\bar{\mathcal{S}}$ and $\mathcal{U}$ are some finite sets. In an IOHMM (536317, ), the state $\bar{S}_{t}$ is hidden and not directly observable, and only an output $\{y_{t}\in\mathcal{Y}\}_{t\in\mathbb{N}}$ is measured, which is a realization from an unknown probability distribution $\mathbb{P}(y_{t}|S_{t},u_{t})$. Here, $\mathcal{Y}$ is some finite set. The objective of the IOHMM is to infer the hidden state $\bar{S}_{t}$ and learn its state transition matrix and $\mathbb{P}(y_{t}|S_{t},u_{t})$ from observables $y_{t}$ and $u_{t}$.
@@ 3.2. Trust-modulated Human Behavior Model
The human behavior model defines the trust evolution and the action selection choice of the human, defining the probability of the human relying on the robot ($o^{+}$) whenever the robot does not seek assistance ($a^{+}$). It is assumed the human’s decision to rely on the robot when it is attempting to collect autonomously is dependent on the complexity of the trial $C_{t}$, and a hidden state which we refer to as the human trust $T_{t}$. We assume that the hidden state of trust can be high $T^{H}$ or low $T^{L}$.
Building upon existing models in the literature (akash2017dynamic, ), we let the trust level $T_{t+1}$ at the end of trial $t$ is influenced by the trust $T_{t}$ at the beginning of the trial, and the experience $E_{t+1}$ at the end of trial $t$ caused by the robot action $a_{t}$ at the complexity $C_{t}$, and the human action $o_{t}$, during trial $t$. Fig. 3 shows the IOHMM-based dynamic human behavior model, which includes the interconnection and influences of the hidden state, inputs, and observable human actions. Training the IOHMM using data sequences results in finding the probability values for trust dynamics $\mathbb{P}(T_{t+1}|T_{t},E_{t+1},C_{t},a_{t})$ and human action selection $\mathbb{P}(o_{t}|T_{t},a_{t},C_{t})$
Figure 3. IOHMM-based dynamic human behavioral model, where the human action $o_{t}$ is modulated by the hidden trust dynamics $T_{t}$
@@ 3.3. POMDP-based Optimal Assistance-Seeking Policy
$\mathbb{P}(S_{t+1}|S_{t},a_{t})=\mathbb{P}(T_{t+1}|E_{t+1},T_{t},C_{t},a_{t})\times\mathbb{P}(C_{t+1})\mathbb{P}(E_{t+1}|T_{t},C_{t},a_{t}),$
@@
$\mathbb{P}(E_{t+1}=E^{+}$ $|o_{t},C_{t},a_{t})=\begin{cases}p_{L}^{\textup{suc}},&\text{if }(a_{t},C_{t},o_{t})=(a^{+},C^{L},o^{+}),\\ p_{H}^{\textup{suc}},&\text{if }(a_{t},C_{t},o_{t})=(a^{+},C^{H},o^{+}),\\ 0,&\text{if }(a_{t},C_{t},o_{t})=(a^{-},*,o^{-}),\\ 0,&\text{if }(a_{t},C_{t},o_{t})=(a^{-},C^{L},*),\\ 1,&\text{if }(a_{t},C_{t},o_{t})=(a^{-},C^{H},*).\end{cases}$
@@
(2) $\begin{split}R(S,a^{+})&=\mathcal{R}^{a^{+}}_{o^{+},E^{+}}\mathbb{P}(E^{+}|o^{+},C,a^{+})\mathbb{P}(o^{+}|T,C,a^{+})+\mathcal{R}^{a^{+}}_{o^{+},E^{-}}\mathbb{P}(E^{-}|o^{+},C,a^{+})\mathbb{P}(o^{+}|T,C,a^{+})+\mathcal{R}^{a^{+}}_{o^{-},E^{+}}\mathbb{P}(o^{-}|T,C,a^{+}),\\ R(S,a^{-})&=\mathcal{R}^{a^{-}}_{*,*},\end{split}$

### v2

# Trust-Aware Assistance-Seeking in Human-Supervised Autonomy

@@ Abstract.
Our goal is to model and experimentally assess trust evolution to predict future beliefs and behaviors of human-robot teams in dynamic environments. Research suggests that maintaining an appropriate level of trust among team members in a human-robot team is vital for team success. Indeed, research suggests that trust is a multidimensional and latent entity that relates to past experiences and future actions in a complex manner. Employing a human-robot collaborative task, we model human behavior using Input-Output Hidden Markov Model (IOHMM) and design an optimal assistance-seeking strategy for the robot using a Partially Observable Markov Decision Process (POMDP) framework. In this task, the human partner supervises an autonomous mobile manipulator collecting objects in an environment. The supervisor’s responsibility is to ensure that the robot safely executes its task. The robot can either choose to attempt to autonomously collect the object or seek human assistance. The human supervisor actively monitors the robot’s activities, offering assistance upon request, and intervening if they perceive that the robot may fail. In this setting, human trust is the hidden state, and the primary objective is to optimize team performance. We conduct two sets of human-robot interaction experiments. The data from the first experiment are used to estimate POMDP parameters, which are used to compute an optimal assistance-seeking policy evaluated in the second experiment. The estimated POMDP parameters reveal that, for most participants, human intervention is more probable when trust is low, particularly in high-complexity tasks. Notably, our estimates suggest that the robot’s action of asking for human assistance in high-complexity tasks can positively impact human trust. Our experimental results show that the proposed trust-aware policy results in a better performance as compared to an optimal trust-agnostic policy. By comparing model estimates of human trust, obtained using only behavioral data, with the collected self-reported trust values, we show that model estimates are isomorphic to self-reported responses.
human-robot interaction, trust modeling, POMDP, IOHMM, behavioral modeling
$c$s: Human-centered computing HCI theory, concepts and modelsccs: Human-centered computing Collaborative interactionccs: Computing methodologies Partially-observable Markov decision processesccs: Computing methodologies Sequential decision makingccs: Computer systems organization Robotics
@@ 1. Introduction
Autonomous systems are becoming increasingly prevalent and are being deployed and used in various sectors such as healthcare, agriculture, military, and transportation. Despite the increasing capabilities of autonomous systems, these systems still require human supervision and assistance when working in complex and uncertain environments(Selma2017, ; akash2020human, ; peterssupervisor2015, ; gupta2024structural, ; Boguslavskii2025, ). However, the increased capabilities of the robots are leading to novel landscapes of human-robot interaction (HRI) wherein the robots are deployed in roles requiring complex social behavior with human partners.
@@
The classical approach to studying the impact of robot reliability, transparency, and human workload on human trust treats trust as a static parameter (hoff2015trust, ). However, recent studies have focused on understanding the dynamic evolution of human-automation trust, focusing on factors such as the quality of robot performance, transparency of the robot’s operation, and the human agent’s attitude and characteristics, such as forgetfulness and self-confidence, on the trust dynamics (lee1992trust, ; lee1994trust, ; manzey2012human, ; desai2013impact, ; akash2017dynamic, ; yang2017evaluating, ; Wang2015DynamicRS, ; chen2020trust, ; akash2020human, ; williams2023computational, ). Hancock et al. (hancock2011meta, ) present a comprehensive review of various factors affecting human trust and categorize them into three major groups: robot-related, human-related, and environment-related factors. They discuss that human trust is most influenced by robot-related factors, which include robot reliability, performance, design, and abilities. This is followed by the environment-related factors, which include task complexity, type, and physical environment. Among the three main categories, human-related factors such as demographics, expertise, and personality have the least influence on trust in human-robot interaction.
@@
Deterministic linear models use a linear time-invariant dynamical system, where the states are the current trust level, cumulative trust level, and expectation bias (akash2017dynamic, ; yang2017evaluating, ; Wang2015DynamicRS, ). The input to these linear models includes human experience, which is determined by the performance and reliability of the robot. Variants of these models have been introduced to capture mutual trust between humans and robots (rahman2018mutual, ; Wang2014, ). These models have been used in several contexts, including exploring the impact of information transparency in reconnaissance missions (yang2017evaluating, ), investigating human reliance on driving assistance systems (akash2017dynamic, ), examining subtask allocation in collaborative assembly tasks (rahman2018mutual, ), and addressing scheduling challenges in teleoperation for underwater robotic navigation (Wang2014, ).
Stochastic models of trust often take one of two forms. Stochastic linear models for the evolution of trust (azevedo2021real, ) rely on linear time-invariant dynamical systems with Gaussian measurement noise. In contrast, other stochastic models treat trust as a categorical latent variable and estimate its distribution based on observable factors like robot performance and human actions (OPTIMO, ; liao2024trust, ). An example of this class of models is the Input-Output Hidden Markov Model (IOHMM) with categorical states, which have been previously used to model human trust (akash2020human, ; DBLP:journals/corr/abs-2009-11890, ; williams2023computational, ). These models are particularly suitable for capturing interpretable internal states (e.g., “High Trust” or “Low Trust”). In addition, unlike linear Gaussian systems, which typically rely on richer or continuous signals such as self-reported trust on Likert scales (chen2020trust, ), gaze duration, or attention time (azevedo2021real, ), or the presence of a secondary task, the IOHMM provide flexibility for handling discrete observations such as human decisions (e.g., Rely or Intervene, Yes or No, Agree or Disagree) in both training and prediction. A comprehensive survey of different trust models is presented in (wang2023human, ). To optimize robot decision-making under uncertainty about internal human states, Partially Observable Markov Decision Process (POMDP) framework has been employed with trust as latent state influenced by robot actions, human responses, and contextual factors (chen2020trust, ; zahedi2023trust, ; akash2020human, ; 10.5555/2906831.2906852, ; DBLP:journals/corr/abs-2009-11890, ). It has been used in applications including scheduling in human-supervised robotics to reduce human interventions (chen2020trust, ) and monitoring load (zahedi2023trust, ), as well as generating optimal recommendations and explanations based on inferred human state (akash2020human, ; 10.5555/2906831.2906852, ). In this work, we combine IOHMM-based trust modeling with a POMDP-based decision-making framework.
@@
In addition to a dynamic trust model, real-time estimation of trust is vital for the design of trust-based policies that are robust to modeling uncertainties. Traditionally, human trust is assessed through self-reporting, such as surveys (schaefer2016measuring, ; lee1992trust, ; muir1989operators, ). However, a limitation of this approach is its impracticality for frequent trust survey completion during the experiment. Psycho-physiological measurements(gebru2022review, ) have also been used to infer the trust state; however, the sensing devices required to collect these measurements, e.g., an EEG head cap, limit their applicability. For a more detailed review of trust measurements, refer to (gebru2022review, ). These challenges have prompted researchers to rely purely on context and behavioral data, such as human actions, to estimate trust. This approach has been implemented and tested in several contexts such as reconnaissance mission (akash2020human, ), autonomous driving (azevedo2021real, ; DBLP:journals/corr/abs-2009-11890, ), and supervision(OPTIMO, ). We focus on similar methodology for estimating human trust: subjective measures and behavioral data.
@@
In this paper, in the context of human-supervised object collection tasks, we design and validate trust-aware policies for an autonomous system seeking assistance from a human supervisor. In contrast to existing works, our focus is on a scenario where the robot can seek assistance from the human supervisor, preventing a loss of trust and possibly repairing it. The major contributions of this work are the following. First, we develop a novel experiment to explore how a robot’s requests for assistance influence human trust. Second, we use an Input-Output Hidden Markov Model (IOHMM) (536317, ) to capture the evolution of the hidden human trust state. In our formulation, the inputs are the robot’s action, environmental complexity, and the outcome or experience in the previous trial, characterized by the success or failure of the robot in accomplishing its goals, and the output is the human action. The present research sought to develop a model of trust in human-robot teams. It differs from previous modeling efforts in 3 different ways. First we investigated trust evolution in a human-robot collaborative task where we evaluated the human’s intervention in the robot’s actions and we sought to understand how human trust is affected if the robot asked for assistance. Using data from multiple participants across several trials, we estimate the parameters of an IOHMM model to explore how assistance-seeking, environmental complexity and prior experience influence human trust and action. The trust estimates with the estimated IOHMM parameters are compared with self-reported survey data. We then combine the IOHMM-based trust modeling with a POMDP-based decision-making framework to design an optimal assistance-seeking policy, comparing its effectiveness with a baseline trust-agnostic policy in further experiments. While IOHMMs and POMDPs have previously been used to model trust and design robot policy, to the best of our knowledge, they have not been applied in the context of assistance-seeking behavior within human-robot collaboration. Our formulation captures how a robot’s choices to seek or avoid help dynamically influence human trust over time.
@@
The paper is structured as follows. Section 2 defines the collaborative object collection task. Section 3 reviews key ideas of IOHMM and POMDP and discusses the design of the optimal trust-aware assistance-seeking policy. The estimated human behavioral model and its implications are described in Section 4. Using the human-in-the-loop experiment, the trust-aware and trust-agnostic policies are compared in Section 5, Section 6 provides an elaborated discussion related to our results, we conclude in Section 7, and Section 8 presents supplementary materials in the Appendix.
@@ 2. Human-Supervised Robotic Object Collection Experiment
Figure 1. Experimental Setup. The top figure shows the world view of the environment available to the human supervisor. The bottom left and middle figures show the local view and end-effector view used by humans during teleoperation. The bottom right figure shows the operation status. The setup uses ROS-Gazebo (Gazebo, ) and resources available in (downs2022google, ; ROSMoMa, ).:
@@
The objects to be collected are placed in two different types of environments: low complexity $C^{L}$ or high complexity $C^{H}$. The complexity of the environment is determined based on the presence of obstacles, obstructing the direct path between the manipulator and the object to be collected. In our experiment, at the beginning of each trial, the complexity of trial is randomly chosen following an independent and identically distributed (i.i.d.) Bernoulli random variable, with $p_{C^{H}}$ representing the probability of the complexity being high.
The outcome of each trial can either be a success or failure. Specifically, autonomous object collection is classified as a failure if the robot collides with its surroundings and/or fails to deposit the object in the bin safely. On the other hand, an autonomous collection is deemed a success if the robot safely collects and deposits the object. In our experiment, the success of autonomous collection in low complexity and high complexity, are mutually independent i.i.d Bernoulli variables with success probabilities $p_{L}^{\textup{suc}}$ and $p_{H}^{\textup{suc}}$, respectively.
@@ 3. Trust-aware Optimal Assistance-seeking
Figure 2. Collaborative object collection setup. The human behavioral model with trust as a hidden state yields the likelihood of human intervention and is used by the robot to design an optimal policy in a POMDP setting. :
@@ 3.1. Background: IOHMM and POMDP
Consider a Markov chain with state $\{\bar{S}_{t}\in\bar{\mathcal{S}}\}_{t\in\mathbb{N}}$ with an input $\{u_{t}\in\mathcal{U}\}_{t\in\mathbb{N}}$, i.e., $\bar{S}_{t}$ is a random process such that the probability distribution of $\bar{S}_{t+1}$ is completely determined by $\bar{S}_{t}$ and $u_{t+1}$. Here, $\bar{\mathcal{S}}$ and $\mathcal{U}$ are some finite sets. In an IOHMM (536317, ), the state $\bar{S}_{t}$ is hidden and not directly observed, and only an output $\{y_{t}\in\mathcal{Y}\}_{t\in\mathbb{N}}$ is measured, which is a realization from an unknown probability distribution $\mathbb{P}(y_{t}|\bar{S}_{t},u_{t})$. Here, $\mathcal{Y}$ is some finite set. The objective of the IOHMM is to infer the hidden state $\bar{S}_{t}$ and learn its state transition matrix and $\mathbb{P}(\bar{S}_{t+1}|\bar{S}_{t},u_{t+1})$ from $y_{t}$ and $u_{t}$.
@@ 3.2. Trust-modulated Human Behavior Model
The human behavior model defines the trust evolution and the action selection choice of the human, defining the probability of the human relying on the robot ($o^{+}$) whenever the robot does not seek assistance ($a^{+}$). It is assumed that the human’s decision to rely on the robot when it is attempting to collect autonomously is dependent on the complexity of the trial $C_{t}$, and a hidden state which we refer to as the human trust $T_{t}$. We assume that the hidden trust state can be high $T^{H}$ or low $T^{L}$.
Building upon existing models in the literature (akash2017dynamic, ), we let the trust level $T_{t+1}$ at the end of trial $t$ as being influenced by the trust $T_{t}$ at the beginning of the trial, and the experience $E_{t+1}$ at the end of trial $t$ caused by the robot action $a_{t}$ at the complexity $C_{t}$, and the human action $o_{t}$, during trial $t$. Fig. 3 shows the IOHMM-based dynamic human behavior model, which includes the interconnection and influences of the hidden state, inputs, and observable human actions. Training the IOHMM using data sequences results in finding the probability values for trust dynamics $\mathbb{P}(T_{t+1}|T_{t},E_{t+1},C_{t},a_{t})$ and human action selection $\mathbb{P}(o_{t}|T_{t},a_{t},C_{t})$
Figure 3. IOHMM-based dynamic human behavioral model, where the human action $o_{t}$ is modulated by the hidden trust dynamics $T_{t}$:
@@ 3.3. POMDP-based Optimal Assistance-Seeking Policy
$\mathbb{P}(S_{t+1}|S_{t},a_{t})=\mathbb{P}(T_{t+1}|E_{t+1},T_{t},C_{t},a_{t})\mathbb{P}(C_{t+1})\mathbb{P}(E_{t+1}|T_{t},C_{t},a_{t}),$
@@
$\mathbb{P}(E_{t+1}=E^{+}$ $|o_{t},C_{t},a_{t})=\begin{cases}p_{L}^{\textup{suc}},&\text{if }(a_{t},C_{t},o_{t})=(a^{+},C^{L},o^{+}),\\ p_{H}^{\textup{suc}},&\text{if }(a_{t},C_{t},o_{t})=(a^{+},C^{H},o^{+}),\\ 0,&\text{if }(a_{t},C_{t},o_{t})=(a^{+},*,o^{-}),\\ 0,&\text{if }(a_{t},C_{t},o_{t})=(a^{-},C^{L},*),\\ 1,&\text{if }(a_{t},C_{t},o_{t})=(a^{-},C^{H},*).\end{cases}$
@@
[…]